When 4.82 g of Ca(s) is added to 273.4 g of water in a container open to the atmosphere, the reaction shown below occurs. The temperature of the solution rises from 21.3 oC to 43.2 oC. What is LaTeX: \DeltaΔH for the reaction if the specific heat of the solution is 4.29J/g oC. (enter the answer with 1 decimal place and no unit, the understood unit is kJ) Ca(s) + 2 H2O(l) ⟶ Ca(OH)2(aq) + H2(g)
Answer #1

The solution of the above question is given below with explanation.

The heat released or absorbed in a reaction can be given by using the formula,

Dleta(H) =m C Delta(T)

where Delta(H)= Heat of the solution

m= mass of the solution

C= specific heat of the solution

And Delta(T),= Final temperature-Initial temperature

Here m= (4.82+ 273..4)g= 278.22 g

C= 4.29 J/g°C

Delta(T) = (43.2-21.3)°C = 21.9 °C

Using the value to find Delta(H)

We get,

Delta (H)= 278.22 x 4.29 x 21.9

Delta (H) = 26139.0 J

= 26.1kJ, is the required answer.

Thanks and regards!

If you have any questions or doubt regarding above you may ask in the comment box.